# Simulating first-order phase transition during inflation

### v1

@@ I Introduction
One highly significant case is the coupling of the inflaton field to a field undergoing a FoPT Ashoorioon (2015); Jiang et al. (2017); Wang et al. (2019); An et al. (2022a). This coupling provides an “graceful exit” mechanism for inflation: the inflaton field rolls slowly during expansion, while the phase-transition field is initially trapped in a metastable vacuum. As the inflaton field evolves, it dynamically adjusts the height of the potential barrier for the phase-transition field. Eventually, the nucleation rate from the false vacuum to the true vacuum becomes large enough for the true vacuum bubbles to rapidly nucleate, expand, and collide against the background exponential expansion, releasing vacuum energy to achieve reheating and end inflation. However, the usual percolation condition with the false-vacuum fraction dropping below $1/e\approx 37\%$ could be misleading during the vacuum-dominated era, as the decreasing fraction in false vacuum might not drop fast enough to compensate for the volume increase in the exponentially expanding space Ellis et al. (2019).
To overcome this problem, we first propose in Sec. II a grand-unification-theory (GUT)-scale FoPT during the Starobinsky inflation Starobinsky (1980), where the height of the phase-transition barrier drops exponentially fast with the field excursion of the inflaton. This ensures no bubble nucleations during most of the inflationary period until the very end of the inflation. Then in Sec. III, we used 3D lattice numerical simulation techniques to simulate this two-field inflation process and obtain the corresponding gravitational wave (GW) energy density spectrum. This spectrum verifies the unique oscillation features in the high-frequency range found in Refs. An et al. (2022a, b, 2024b, 2024a), providing a distinct physical signal for future GW detection. We conclude in Sec. IV with some discussions. Throughout Sec. III, we adopt units where $c=8\pi G=1$.
@@ II.1 A GUT FoPT during Starobinsky inflation
where the total potential $U(\chi,\phi)=W(\chi)+V(\phi)$ consists of a plateau potential $W(\chi)$ as in the Starobinksy inflation Starobinsky (1980); Kehagias et al. (2014); Whitt (1984); Maeda (1988); Gorbunov and Panin (2011), and a FoPT potential $V(\phi)$ with the height of potential barrier controlled by the parameter $a$,
@@ III.2 Simulation results
We then turn to the properties of the GW spectrum in this model. Fig. 6 presents the GW power spectra at four representative moments. As bubbles nucleate and collide within the $\phi$ field, GWs begin to be generated. Here, we define the phase transition time as the moment when the bubble volume fraction reaches 0.7 Ellis et al. (2019); Liu et al. (2022), which is indicated by the vertical line. The signal eventually stabilizes around $\tau/R^{*}=1.95$. Notably, in the high-frequency regime where $\text{log}_{10}[kR^{*}]>1$, the power spectrum exhibits a distinct oscillatory pattern. This numerical finding is in agreement with the theoretical predictions for inflation-triggered FoPTs An et al. (2022b). During the early stage, the universe is dominated by vacuum energy, and the $\mathcal{H}$ remains relatively stable with little variation. In this background, the GW modes generated by the phase transition are stretched beyond the Hubble horizon due to the rapid expansion and become “freeze out”. Because modes with different wavelengths exit the horizon at different times, they accumulate distinct phases during the subsequent evolution. Upon re-entering the horizon, the coherent superposition of these modes leads to the characteristic oscillatory structure observed in the power spectrum.
IV Consclusions and discussions
@@
Several aspects merit further improvements and investigations in future works: First, numerical simulations of the whole inflationary era, not just near the end of inflation, are necessary for a rigorous investigation of this two-field dynamics; Second, successive bubble nucleations near the end of inflation, not just at one characteristic moment (the moment that nucleated bubbles collide within one e-folding number), are necessary for a complete survey of bubble collisions, including the collisions between large/rare and small/numerous bubble collisions; Third, particle production might occur by bubble collisions during inflation Shakya (2025); Mansour and Shakya (2025); Inomata et al. (2025), and even if bubbles cannot collide during inflation for some FoPT model buildings, it is still of great observational interest to study the primordial black hole formation Deng and Vilenkin (2017); Deng (2020) and isocurvature perturbations Buckley et al. (2024).

### v2

@@ I Introduction
One highly significant case is the coupling of the inflaton field to a field undergoing a FoPT Ashoorioon (2015); Jiang et al. (2017); Wang et al. (2019); An et al. (2022a). This coupling provides a “graceful exit” mechanism for inflation: the inflaton field rolls slowly during expansion, while the phase-transition field is initially trapped in a metastable vacuum. As the inflaton field evolves, it dynamically adjusts the height of the potential barrier for the phase-transition field. Eventually, the nucleation rate from the false vacuum to the true vacuum becomes large enough for the true vacuum bubbles to rapidly nucleate, expand, and collide against the background exponential expansion, releasing vacuum energy to achieve reheating and end inflation. However, the usual percolation condition with the false-vacuum fraction dropping below $1/e\approx 37\%$ could be misleading during the vacuum-dominated era, as the decreasing fraction in false vacuum might not drop fast enough to compensate for the volume increase in the exponentially expanding space Ellis et al. (2019).
To overcome this problem, we first propose in Sec. II a grand-unification-theory (GUT)-scale FoPT Hu and Zhou (2025) during the Starobinsky inflation Starobinsky (1980), where the height of the phase-transition barrier drops exponentially fast with the field excursion of the inflaton. This ensures no bubble nucleations during most of the inflationary period until the very end of the inflation. Then in Sec. III, we used 3D lattice numerical simulation techniques to simulate this two-field inflation process and obtain the corresponding gravitational wave (GW) energy density spectrum. This spectrum verifies the unique oscillation features in the high-frequency range found in Refs. An et al. (2022a, b, 2024b, 2024a); Hu and Zhou (2025), providing a distinct physical signal for future GW detection. We conclude in Sec. IV with some discussions. Throughout Sec. III, we adopt units where $c=8\pi G=1$.
@@ II.1 A GUT FoPT during Starobinsky inflation
where $M_{\mathrm{Pl}}$ is the reduced Planck mass, the total potential $U(\chi,\phi)=W(\chi)+V(\phi)$ consists of a plateau potential $W(\chi)$ as in the Starobinksy inflation Starobinsky (1980); Kehagias et al. (2014); Whitt (1984); Maeda (1988); Gorbunov and Panin (2011), and a FoPT potential $V(\phi)$ with the height of potential barrier controlled by the parameter $a$,
@@ III.2 Simulation results
We then turn to the properties of the GW spectrum in this model. Fig. 6 presents the GW power spectra at four representative moments. As bubbles nucleate and collide within the $\phi$ field, GWs begin to be generated. Here, we define the phase transition time as the moment when the bubble volume fraction reaches 0.7 Ellis et al. (2019); Liu et al. (2022), which is indicated by the vertical line. The signal eventually stabilizes around $\tau/R^{*}=1.95$. Notably, in the high-frequency regime where $\text{log}_{10}[kR^{*}]>1$, the power spectrum exhibits a distinct oscillatory pattern. This numerical finding is in agreement with the theoretical predictions for inflation-triggered FoPTs An et al. (2022b). During the early stage, the universe is dominated by vacuum energy, and $\mathcal{H}$ remains relatively stable with little variation. In this background, the GW modes generated by the phase transition are stretched beyond the Hubble horizon due to the rapid expansion and become “freeze out”. Because modes with different wavelengths exit the horizon at different times, they accumulate distinct phases during the subsequent evolution. Upon re-entering the horizon, the coherent superposition of these modes leads to the characteristic oscillatory structure observed in the power spectrum.
Figure 7: GW spectra from GUT phase transition on different scenarios $N^{*}=0.9,13,17,22$. The sensitivity curves data comes from Hu and Wu (2017); Ruan et al. (2020); Amaro-Seoane and others (2017); Baker and others (2019); Crowder and Cornish (2005); Corbin and Cornish (2006); Harry et al. (2006); Seto et al. (2001); Kawamura and others (2006); Yagi and Seto (2011); Isoyama et al. (2018); Lentati and others (2015); Shannon and others (2015); Arzoumanian and others (2018); Sesana and others (2021); Carilli and Rawlings (2004); Aasi and others (2015); Reitze and others (2019); Punturo and others (2010); Cyburt et al. (2016); Ejlli et al. (2019); Domcke and Garcia-Cely (2021); Ito et al. (2024); Dandoy et al. (2024).
After being redshifted to the present epoch, the peak frequency of the gravitational wave (GW) signal is given by $f_{p}=1.6\times 10^{-5}\frac{\beta}{H_{*}}\frac{T_{*}}{100}(\frac{g_{*}}{100})^{1/6}\text{ Hz}$, with a corresponding spectral amplitude of $\Omega(f_{p})=1.67\times 10^{-5}(\frac{100}{g_{*}(T_{*})})^{1/3}(\frac{\alpha}{1+\alpha})^{2}(\frac{H_{*}}{\beta})^{2}$. The resulting redshifted GW spectra are presented in Fig. 7, alongside the sensitivity curves of various GW detectors. Furthermore, by employing the model from Ref. Hu and Zhou (2025), When the phase transition is treated as a transient or instantaneous source occurring at the inflationary moment $N^{*}$, its peak frequency scales as $f_{p}\propto e^{-N^{*}}$. we can achieve an earlier phase transition with $N^{*}=13,17,\text{ and }22$. Consequently, the generated gravitational waves undergo a more pronounced redshift, shifting the signal into lower-frequency bands accessible to current and future detectors.
IV Conclusions and discussions
@@
Several aspects merit further improvements and investigations in future works: First, numerical simulations of the whole inflationary era, not just near the end of inflation, are necessary for a rigorous investigation of this two-field dynamics; Second, successive bubble nucleations near the end of inflation, not just at one characteristic moment (the moment that nucleated bubbles collide within one e-folding number), are necessary for a complete survey of bubble collisions, including the collisions between large/rare and small/numerous bubble collisions; Third, particle production might occur by bubble collisions during inflation Shakya (2025); Mansour and Shakya (2025); Inomata et al. (2025), and even if bubbles cannot collide during inflation for some FoPT model buildings, it is still of great observational interest to study the primordial black hole formation Deng and Vilenkin (2017); Deng (2020); Konoplich et al. (1999); Khlopov et al. (2000, 2000); Dymnikova et al. (2000); Khlopov (2024) and isocurvature perturbations Buckley et al. (2024).